OpenStudy (mopsy510):

Are you able to graph these quadratic equations by finding the x and y intercept and the vertex? x^2+5x+6=y x^2-5x=y 3x^2+12x+12

OpenStudy (anonymous):

x^2+5x+6=y x^2-5x=y x^2+5x+6=x^2-5x , 10x=-6 , x=-0,6 (-0,6)^2+5(-0,6)+6=y , y = 3,36 That's the location ain't it?